Jbaena1:

The organic garden at a restaurant is rectangular in shape and has a perimeter of 84 feet. If the length l of the garden is 24 feet longer than the width w of the garden, what is the area of the garden, in square feet? Possible Answers: A. 9 B. 33 C. 84 D. 297 E. I don't know.

Jbaena1:

Its sounds so simple but the one guy who was trying to teach me made it over complicated.

Vocaloid:

Perimeter of a rectangle is P = 2L + 2W “Length is 24 more than the width” so L = W + 24 That equals sign means that we can replace L with W + 24 whenever L shows up So going back to the original equation for perimeter P = 2L + 2W = 2(W + 24) + 2W = 84 since perimeter is already given Solve for W Now, you can’t just stop there because it wants the area not the perimeter So you must then find L and then find area = L*W